OpenStudy (athenaswisegirl):

Find the area bounded by the curves x = y^2 – 4y and x = 2y – y^2. Your work must include an integral in one variable

OpenStudy (sshayer):

x=y^2-4y add 4 to both sides x+4=y^2-4y+4 (y-2)^2=(x+4) so vertex of this parabola is (-4,2) x=2y-y^2 y^2-2y=-x y^2-2y+1=-x+1 (y-1)^2=-(x-1) vertex of this parabola is (1,1)